Exercise 4:

Consider a financial institution with • Assets of 120 AUD
• Uninsured deposits of 50 AUD • Insured deposits 50 AUD

• Equity 20 AUD

a) The financial institution faces a loss of 10 AUD. Calculate the asset value, the value to the previous uninsured deposit holders, insured deposit holders, and equity holders. Does the financial institution run into bankruptcy? Does the deposit insurance have to pay?

b) The financial institution faces a loss of 30 AUD. Calculate the asset value, the value to the previous uninsured deposit holders, insured deposit holders, and equity holders. Does the financial institution run into bankruptcy? Does the deposit insurance have to pay?

c) The financial institution faces a loss of 80 AUD. Calculate the asset value, the value to the previous uninsured deposit holders, insured deposit holders, and equity holders. Does the financial institution run into bankruptcy? Does the deposit insurance have to pay?
Answer #1

The sequence will be as follows:

1. Losses will first wipe out the equity

2. It will then wipe out the uninsured deposits

The insured deposits will remain the same throughout and the deposit insurance company will have to absorb the losses.

Please see the tables below:

Part (a)

Assets Value (AUD) Liabilities & Equity Value (AUD)
Assets 120 - 10 = 110 Uninsured Deposits 50
Insured Deposits 50
Equity 20 - 10 = 10
Total 110 Total 110

Does the financial institution run into bankruptcy?

Since the value of equity > 0, the financial institution has yet not run into bankruptcy.

Does the deposit insurance have to pay?

Since, the insured deposits are still intact, the deposit insurance need not pay.

Part (b)

Assets Value (AUD) Liabilities & Equity Value (AUD)
Assets 120 - 30 = 90 Uninsured Deposits 50 - 10 = 40
Insured Deposits 50
Equity 20 - 20 = 0
Total 90 Total 90

Does the financial institution run into bankruptcy?

Since the value of equity = 0, the financial institution has run into bankruptcy.

Does the deposit insurance have to pay?

Since, the insured deposits are still intact, the deposit insurance need not pay.

Part (c)

Assets Value (AUD) Liabilities & Equity Value (AUD)
Assets 120 - 80 = 40 Uninsured Deposits 50 - 50 = 0
Insured Deposits 50
Equity 20 - 20 = 0
Total 40 Total 50

Does the financial institution run into bankruptcy?

Since the implied value of equity = 40 - 50 = -10 < 0, the financial institution has run into bankruptcy.

Does the deposit insurance have to pay?

Based on its assets, the firm can pay only AUD 40 to its insured depositor. Hence the deposit insurance has to pay an amount = AUD 50 - 40 = AUD 10